OpenStudy (anonymous):

a block of ice in the shape of a cube originally has a volume of 1000 cm^3. as it melts each of the edges is dereasing at a rate of 1cm/hr. At what rate is the surface area of the cube decreasing when its volume is 27 cm^3

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

please..

OpenStudy (anonymous):

the surface area of a cube is 6x^2, the derivative is 12x so the rateof change when volume is 27( x= 3 ) is 12*3 = 36 cm^2 per hour I think

OpenStudy (dumbcow):

Let x be length of edge of cube SA = 6x^2 dSA/dx = 12x dx/dt = -1 dSA/dt = dx/dt * dSA/dx dSA/dt = -12x Now Volume = x^3 x^3 = 27 x = 3 Therefore, dSA/dt = -12*3 = -36

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

its cm^2/hr right?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

cool man

OpenStudy (ash2326):

let the edge of the cube be a da/dt=-1cm/hr originally the volume was 1000 cm^3 so a^3=1000 a=10 cm initially surface area s=6a^2 we have to find ds/dt when v=27 cm^3 or a=3 cm differentiate s with respect to t ds/dt=12a da/dt ds/dt=12*3*-1 ds/dt=-36 =-36 cm^2/hr